OpenStudy (anonymous):

What are the odds of rolling two number cubes and getting a sum of seven? A.) 1/6 B.) 1/7 C.) 5/31 D.) 1/5

OpenStudy (anonymous):

in my time we called these things "dice"

OpenStudy (anonymous):

there are 36 possible rolls

OpenStudy (anonymous):

I know right? I don't see why they can't just say dice instead of number cubes.

OpenStudy (phi):

number cube!? gack!

OpenStudy (anonymous):

of those 36 possible rolls, you have six ways to get a 7, listed here (1,6) (2,5) (3,4) (4,3) (5,2) (6,1)

OpenStudy (anonymous):

Lol seriously, though. @phi

OpenStudy (anonymous):

36/7 is 5 point something on my calculater so the answer is c :) hope this helped

OpenStudy (anonymous):

the ratio is therefore \(\frac{6}{36}=\frac{1}{6}\)

OpenStudy (anonymous):

Two different answers...ahhh

OpenStudy (phi):

yes, but only sat has a 100 score

OpenStudy (anonymous):

type it into your calculator and it'll show you

OpenStudy (anonymous):

first of all one thing should be clear: a probability is always a number between zero and one

OpenStudy (anonymous):

howd u get 6 36 in the firt place??? jw

OpenStudy (anonymous):

in this case it is the ratio of the number of favorable events divided by the total number of events, since they are all equally likely

OpenStudy (anonymous):

@callie2240 one "number cube" has 6 faces, the other also has 6 faces total number of possible outcomes, by the counting principle, is \(6\times 6=36\)

OpenStudy (anonymous):

I have more questions if y'all wanna help :)

OpenStudy (anonymous):

sure and ok, i'm sorry i messed up i didnt think of that one @satellite73